Solve.
The length of a rectangle is 8 in. more than twice its width. If the perimeter of the rectangle is 34 in., find the width of the rectangle.

Answer by checkley71(8403)

L=2W+8
2L+2W=34
2(2W+8)+2W=34
4W+16+2W=34
6W=34-16
6W=18
W=18/6
W=3 ANSWER FOR THE WIDTH.
L=2*3+8
L=6+8
L=14 ANSWER.
PROOF.
2*14+2*3=34
28+6=34
34=34